OpenStudy (amtran_bus):

*Logic* While driving from city A to B, a car got 22 MPG and while returning on the same road, it got 30 MPG.

OpenStudy (amtran_bus):

Which quantity is greater? A: Car's average MPG for both trips MPG B: 26

OpenStudy (amtran_bus):

The answer is 26...But I do not understand.

OpenStudy (amtran_bus):

Like, why are they not equal?

OpenStudy (astrophysics):

(22+30)/2 MPG = 26 MPG

OpenStudy (astrophysics):

They just want you to get the average of going and coming back

OpenStudy (amtran_bus):

Yep! The GRE book says they are not equal though, that quantity A comes out to be about 25.38.

OpenStudy (amtran_bus):

They say the average will be weighted more to the side of 22 for some reason.

OpenStudy (amtran_bus):

They say A--> B is 22*30 = 660 B--> A is 30 *22 = 660 Total is x * 52 = 1320 x is 25.38

OpenStudy (astrophysics):

I don't understand, what is the question?

OpenStudy (amtran_bus):

While driving from city A to B, a car got 22 MPG and while returning on the same road, it got 30 MPG. Which quantity is greater? A: Car's average MPG for both trips MPG B: 26 That is literally it.

zepdrix (zepdrix):

I was thinking of it as Astro was, a simple arithmetic mean of the two rates... but I guess they're doing something else here.. hmm

OpenStudy (amtran_bus):

I have spent awhile trying to figure their solution out but can't. I think it is just poorly worded. But that's what the question and their solution says. I would never get that right on a test.

OpenStudy (amtran_bus):

@Astrophysics do you think it is something we are missing or is it just worded poorly? This is GRE review book.

OpenStudy (astrophysics):

Hey, sorry I'm not really sure :\

OpenStudy (amtran_bus):

It's all good! Thanks for trying!

OpenStudy (astrophysics):

They have some sort of ratio going on, but I don't see the logic

OpenStudy (amtran_bus):

I mean, if A--->B was longer than b-->A, it would make sense.

OpenStudy (astrophysics):

@ganeshie8

satellite73 (satellite73):

the average gas mileage is not the average of the gas mileages

OpenStudy (amtran_bus):

That short and supposingly simple statement just blew me away.

satellite73 (satellite73):

exactly the same as "the average speed is not the average of the speeds" an example with numbers will make it clear

satellite73 (satellite73):

lets say the distance from A to B is 100 miles (not that it makes any difference)

OpenStudy (amtran_bus):

I'm following. Thanks.

OpenStudy (aleah54):

do you still need help

satellite73 (satellite73):

if they got 22 miles per gallon then they used up how many gallons?

OpenStudy (amtran_bus):

Like 4.5, right?

OpenStudy (aleah54):

oh ok if you still don't get it just tag me

satellite73 (satellite73):

just as a fraction, not as a decimal it would be \[\frac{100}{22}\]

OpenStudy (amtran_bus):

Right.

OpenStudy (amtran_bus):

Ok, so if they got 30 MPG it would be 100/30

satellite73 (satellite73):

which is not 4.5, but whatever now coming back, they use up \[\frac{100}{30}\]

satellite73 (satellite73):

how far have they gone? 200 how much did they use? \[\frac{100}{22}+\frac{100}{30}\]

satellite73 (satellite73):

i realize you can reduce these fractions, but no need so what is there average miles per gallon? \[\frac{200}{\frac{100}{22}+\frac{100}{30}}\]

OpenStudy (amtran_bus):

25.38

OpenStudy (amtran_bus):

That makes so much more sense.

satellite73 (satellite73):

that is the total distance 200 divided by the total amount used

OpenStudy (amtran_bus):

Thank you SO MUCH!

satellite73 (satellite73):

now if you want to do it with D instead of 100, you get the same answer

satellite73 (satellite73):

wanna try it?

OpenStudy (astrophysics):

Nice! I'm going to say I didn't get it because it wasn't using a metric system yeah! :)

satellite73 (satellite73):

yw btw wanna try with D?

OpenStudy (amtran_bus):

Well, if we used D, it would give the same answer. I mean, it is just sub for miles.

satellite73 (satellite73):

\[\frac{2D}{\frac{D}{22}+\frac{D}{30}}\] is the start then a bit of algebra, the D will go away

OpenStudy (amtran_bus):

Just multiply by inverse, right?

OpenStudy (amtran_bus):

So many viewing this! We are famous.

satellite73 (satellite73):

add in the denominator might be easiest

satellite73 (satellite73):

\[\frac{2D}{\frac{22D+30D}{22\times 30}}\] is a start

satellite73 (satellite73):

btw this has a name, think it might be "harmonic mean" but i could be wrong

OpenStudy (amtran_bus):

lol it would help if I used the equation feature.

satellite73 (satellite73):

yeah it is replace D by 1 you get \[\frac{2}{\frac{1}{22}+\frac{1}{30}}\] pretty sure this is harmonic mean google

OpenStudy (amtran_bus):

=25.38. Thanks.

OpenStudy (amtran_bus):

https://en.wikipedia.org/wiki/Harmonic_mean

satellite73 (satellite73):

yup harmonic mean it is

OpenStudy (amtran_bus):

That's neat we ran into this in this supposingly simple problem. Thanks of all your help. I learned a lot.

OpenStudy (aleah54):

:)

satellite73 (satellite73):

oh also improtant inequatly arithmetic mean > geometric mean > harmonic mean by > it is really \(\geq\)